### v1

# references.bib

Abstract
The chromatic symmetric $X_{G}$ function is a symmetric function generalization of the chromatic polynomial of a graph, introduced by Stanley (1995). $X_{G}$ is defined by assigning a variable to each color and taking the sum over all proper vertex colorings of $G$ of the product of all variables corresponding to colors used, counted with multiplicity. Stanley (1995) gave an expansion formula for $X_{G}$ in terms of the power sum symmetric functions $p_{\lambda}$ using the principle of inclusion-exclusion. Crew, Pechenik, and Spirkl (2023) defined the Kromatic symmetric function $\overline{X}_{G}$ as a $K$-theoretic analogue of $X_{G}$, constructed in the same way except that each vertex is assigned a nonempty set of colors such that adjacent vertices have nonoverlapping color sets. They defined a $K$-analogue $\overline{p}_{\lambda}$ of the power sum basis and computed the first few coefficients of the $\overline{p}$-expansion of $\overline{X}_{G}$ for some small graphs $G$. They conjectured that the $\overline{p}$-expansion always has integer coefficients and asked whether there is an explicit formula for these coefficients. We give a formula for the $\overline{p}$-expansion of $\overline{X}_{G}$, and prove that the coefficients are indeed always integers.
1 Introduction
The chromatic symmetric function $X_{G}$ was introduced by Stanley in [stanley1995symmetric] as a symmetric function generalization of the chromatic polynomial [birkhoff1913reducibility]. Stanley gave expansion formulas for $X_{G}$ in a few different bases, including the power sum symmetric functions $p_{\lambda}.$ His classic $p$-expansion formula is an alternating sum over all subsets of the edges of $G$, but he also gave an alternative formulation in terms of the broken circuit complex of $G$, which unlike his subset sum expansion does not involve extra cancellation. Bernardi and Nadeau ([bernardi2020combinatorial]) gave an alternative interpretation of the $p$-expansion of $X_{G}$ in terms of the source components of its acyclic orientations.
The Kromatic symmetric function $\overline{X}_{G}$ was defined by Crew, Pechenik, and Spirkl in [kromatic2023] as a $K$-theoretic analogue of $X_{G}.$ The idea of $K$-theory is to deform the cohomology ring of a topological space by introducing an extra parameter $\beta$ (often set to $-1$), so part of the motivation for defining $\overline{X}_{G}$ was to help shed light on possible topological interpretations for $X_{G}.$ Marberg [marberg2023kromatic] gave an alternative construction of $\overline{X}_{G}$ in terms of linearly compact Hopf algebras and showed that it arises as a natural analogue of $X_{G}$ in that setting as well.
One natural question about $\overline{X}_{G}$, posed by the authors of [kromatic2023], is whether there is a $K$-analogue of Stanley’s power sum expansion formula in terms of an appropriate $K$-analogue of the $p$-basis. The authors of [kromatic2023] defined such a $K$-analogue of the $p$-basis, which they called the $\overline{p}$-basis. They gave a table listing the first few coefficients of the $\overline{p}$-expansions of $\overline{X}_{G}$ for several small graphs $G$, which they computed using Sage [sage]. The leading terms (i.e. the terms where $|\lambda|$ is minimal), must always match the corresponding terms from ordinary $p$-expansion of the ordinary chromatic symmetric function $X_{G}.$ (The leading terms are those where $|\lambda|=|V(G)|$ for an unweighted graph, or where $|\lambda|=\omega(G):=\sum_{v\in V(G)}\omega(v)$ for a weighted graph with weight function $\omega:V(G)\to\mathbb{N}$.)
Unlike the $p$-expansion of $X_{G},$ they found that the $\overline{p}$-expansions for $\overline{X}_{G}$ appeared to have infinitely many terms, but the coefficients still seemed to always be integers. For instance, below are first few terms (grouped by $|\lambda|$) of the $\overline{p}$-expansions for the single edge [Uncaptioned image], and for the graph $K_{21}$ consisting of a single edge connecting a vertex of weight 2 and a vertex of weight 1:
$\overline{X}_{K_{21}}$ $=(-\overline{p}_{3}+\overline{p}_{21})+(\overline{p}_{4}-\overline{p}_{31})+(\overline{p}_{41}-\overline{p}_{32})+(-\overline{p}_{6}+\overline{p}_{42}+\overline{p}_{33}-\overline{p}_{321})$
$\ \ \ +(-\overline{p}_{61}+\overline{p}_{421}+\overline{p}_{331})+(\overline{p}_{8}-\overline{p}_{62}+\overline{p}_{332})+(\overline{p}_{81}+\overline{p}_{63}-\overline{p}_{621}-\overline{p}_{333}+\overline{p}_{3321})+\dots$
[…]
$\ \ \ +(6\overline{p}_{5}-8\overline{p}_{41}-3\overline{p}_{32}+2\overline{p}_{311}+2\overline{p}_{221})$
$\ \ \ +(-9\overline{p}_{6}+12\overline{p}_{51}+4\overline{p}_{42}-4\overline{p}_{411}+\overline{p}_{33}-4\overline{p}_{321}-\overline{p}_{222}+\overline{p}_{2211})\dots.$
[…]
The authors of [kromatic2023] asked whether there is an explicit formula for the $\overline{p}$-expansion of $\overline{X}_{G}$. Here, we give such a formula that applies to any weighted graph $G$ with a vertex weight function $\omega:V\to\mathbb{N},$ in terms of its independence polynomial
$I_{(G,\omega)}(t):=\sum_{S\subseteq V(G)\text{ an independent set}}t^{\omega(S)},$
where the weight of a subset $S$ is $\omega(S):=\sum_{v\in S}\omega(v).$ We write $G|_{W}$ to denote the induced subgraph of $G$ with vertex set $W.$
Theorem 1.
$\overline{X}_{(G,\omega)}$ satisfies the formula
$\overline{X}_{(G,\omega)}=\sum_{W\subseteq V(G)}(-1)^{|V(G)|-|W|}\prod_{k\geq 1}(1+\overline{p}_{k})^{a_{W}(k)},$
where the exponents $a_{W}(k)$ are defined to satisfy
$(1+t)^{a_{W}(1)}(1+t^{2})^{a_{W}(2)}(1+t^{3})^{a_{W}(3)}\dots=I_{(G|_{W},\omega)}(t).$
This idea for where the $(1+t^{k})$ terms come from is the factorization $1+\overline{p}_{k}=\prod_{i\geq 1}(1+x_{i}^{k}).$ As a consequence of Theorem 1, we can write down an explicit formula for the coefficients $[\overline{p}_{\lambda}]\overline{X}_{(G,\omega)}$:
Corollary 2.
The coefficient of $\overline{p}_{\lambda}$ in $\overline{X}_{(G,\omega)}$ is given by
$[\overline{p}_{\lambda}]\overline{X}_{(G,\omega)}=\sum_{W\subseteq V(G)}(-1)^{|V(G)|-|W|}\prod_{k=1}^{\ell}\binom{a_{W}(k)}{i_{k}},$
where $\lambda=\ell^{i_{\ell}}\dots 2^{i_{2}}1^{i_{1}}.$
The numbers $a_{W}(k)$ above can be computed recursively as follows:
Proposition 3.
The exponents $a_{W}(k)$ satisfy the recurrence
$a_{W}(k)=[t^{k}]\log(I_{(G|_{W},\omega)}(t))+\sum_{d\mid k}\frac{(-1)^{k/d}d\cdot a_{W}(d)}{k}.$
We also show that all coefficients in the above expansion are integers:
Theorem 4.
For every partition $\lambda,$ the coefficient $[\overline{p}_{\lambda}]\overline{X}_{G}$ is an integer.
Another question one can ask is a question of $\overline{p}$-positivity. For $G$ an unweighted graph, it can be shown that the sign of $[p_{\lambda}]X_{G}$ is $(-1)^{|\lambda|-\ell(\lambda)}$. From checking some examples, the coefficients $[\overline{p}_{\lambda}]\overline{X}_{G}$ seem to follow the same pattern of signs, which suggests the following conjecture:
Conjecture 5.
For $G$ unweighted, the sign of $[\overline{p}_{\lambda}]\overline{X}_{G}$ is $(-1)^{|\lambda|-\ell(\lambda)}.$
Note that Conjecture 5 does not apply to weighted graphs. For instance, in the $\overline{p}$-expansion of $K_{21}$ above, the coefficient for $\lambda=3$ is $-1$ while $|\lambda|-\ell(\lambda)=3-1=2$ is even.
The remainder of this paper is organized as follows: In §2, we define some necessary terminology, in §3, we revisit the examples above to illustrate Theorem 1, and in §4, we prove Theorem 1, Corollary 2, Proposition 3, and Theorem 4.
2 Background
A graph $G$ is a set $V(G)$ of vertices together with a set $E(G)$ of edges, where each edge is an unordered pair of vertices. If $uv$ is an edge, we say that the vertices $u$ and $v$ are adjacent. A weighted graph $(G,\omega)$ is a graph $G$ together with a weight function $\omega:V(G)\to\mathbb{N}$ that assigns a positive integer to each vertex. An independent set or stable set is a subset of $V(G)$ in which no two vertices are adjacent. For $W\subseteq V,$ the induced subgraph $G|_{W}$ with vertex set $v$ is the graph with vertex set $V(G|_{W}):=W$ and edge set $E(G|_{W}):=\{uv\in E(G):u,v\in W\}.$ That is, the edges of $G|_{W}$ are precisely the edges of $G$ with both endpoints in $W.$
A symmetric function $f(x_{1},x_{2},\dots)$ is a function that stays the same under any permutation of the variables, i.e. $f(x_{1},x_{2},\dots)=f(x_{\sigma(1)},x_{\sigma(2)},\dots)$ for any permutation $\sigma$ of $\mathbb{N}.$ A proper coloring of $V(G)$ is a function $\alpha:V(G)\to\mathbb{N}$ that assigns a positive integer valued color to each vertex such that $\alpha(u)\neq\alpha(v)$ whenever $uv\in E(G).$ The chromatic symmetric function is
$X_{G}:=\sum_{\alpha}\prod_{v\in V(G)}x_{\alpha(v)}$
is the sum over all proper set colorings of the product of all the variables corresponding to colors used, counted with multiplicity.
A partition $\lambda=\lambda_{1}\lambda_{2}\dots\lambda_{k}$ is a nonincreasing sequence of positive integers $\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{k},$ and $\lambda_{1},\dots,\lambda_{k}$ are its parts. We can alternatively write $\lambda=\ell^{i_{\ell}}(\ell-1)^{i_{\ell-1}}\dots 2^{i_{2}}1^{i_{1}}$ to denote the partition with $i_{j}$ parts of size $j$ for each $j=1,2,\dots,\ell-1,\ell.$ For a partition $\lambda=\lambda_{1}\lambda_{2}\dots\lambda_{k},$ the power sum symmetric function $p_{\lambda}$ is
$p_{\lambda}:=\prod_{i=1}^{k}\left(\sum_{j\in\mathbb{N}}x_{j}^{\lambda_{i}}\right)=(x_{1}^{\lambda_{1}}+x_{2}^{\lambda_{1}}+\dots)(x_{1}^{\lambda_{2}}+x_{2}^{\lambda_{2}}+\dots)\dots(x_{1}^{\lambda_{k}}+x_{2}^{\lambda_{k}}+\dots).$
Stanley (1995, [stanley1995symmetric]) proved the following $p$-expansion formula for $X_{G}$:
Theorem 6 (Stanley [stanley1995symmetric], Theorem 1.1).
$X_{G}$ satisfies
$X_{G}=\sum_{S\subseteq E(G)}(-1)^{|S|}p_{\lambda(S)},$
where $\lambda(S)$ is the partition whose parts are the sizes of the connected components of the subgraph of $G$ with vertex set $V(G)$ and edge set $S.$
A proper set coloring is a function $\alpha:V(G)\to 2^{\mathbb{N}}\backslash\{\varnothing\}$ that assigns a nonempty set of positive integer valued colors to each vertex such that adjacent vertices have nonoverlapping color sets, i.e. $\alpha(u)\cap\alpha(v)=\varnothing$ whenever $uv\in E(G).$ For a weighted graph $(G,\omega)$, the Kromatic symmetric function is
$\overline{X}_{(G,\omega)}:=\sum_{\alpha}\prod_{v\in V(G)}\prod_{i\in\alpha(v)}x_{i}^{\omega(v)}.$
The authors of [kromatic2023] define the $K$-analogue of the power sum symmetric functions as
$\overline{p}_{\lambda}:=\overline{X}_{\overline{K_{\lambda}}},$
where $\overline{K_{\lambda}}$ is the edgeless graph whose vertex weights are the parts of $\lambda.$
We will use a few more definitions to prove Theorem 4. The monomial symmetric function $m_{\lambda}$ is
$m_{\lambda}:=\sum_{n_{1},\dots,n_{k}}x_{n_{1}}^{\lambda_{1}}\dots x_{n_{k}}^{\lambda_{k}},$
where the sum is taken over all ordered tuples of $k$ distinct indices $n_{1},\dots,n_{k}\in\mathbb{N}$. The augmented monomial symmetric function $\widetilde{m}_{\lambda}$ is
$\widetilde{m}_{\lambda}:=i_{1}!\dots i_{\ell}!\cdot m_{\lambda},$
where $i_{j}$ is the number of parts of $\lambda$ of size $j.$ The $K$-theoretic augmented monomial symmetric functions $\overline{\widetilde{m}}_{\lambda}$ is
$\overline{\widetilde{m}}_{\lambda}:=\overline{X}_{K_{\lambda}},$
where $K_{\lambda}$ is a complete graph whose vertex weights are the parts of $\lambda.$ A stable set cover $C=\{S_{1},\dots,S_{k}\}$ is a set of $k$ independent sets in $G$ such that every vertex is in at least one of $S_{1},\dots,S_{k}.$ We write $\textsf{SSC}(G)$ for the set of all stable set covers of $G$, and $\lambda(C)$ for the partition whose parts are the weights $\omega(S_{1}),\dots,\omega(S_{k}),$ where the weight of a subset is the sum of the weights of its vertices, $\omega(S):=\sum_{v\in S}\omega(v).$ The authors of [kromatic2023] prove the following $\overline{\widetilde{m}}$-expansion formula for $\overline{X}_{G}$:
Theorem 7 (Crew, Pechenik, and Spirkl [kromatic2023], Proposition 3.4).
The $\overline{\widetilde{m}}$-expansion for $\overline{X}_{G}$ is
$\overline{X}_{(G,\omega)}=\sum_{C\in\textsf{\emph{SSC}}(G)}\overline{\widetilde{m}}_{\lambda(C)}.$
3 Examples of Theorem 1
[…]
Example 8.
Let $G=K_{21}$ consist of two connected vertices $v$ and $w,$ with $\omega(v)=1$ and $\omega(w)=2$. Then
$I_{(G,\omega)}(t)=1+t+t^{2},$
since $G$ has three independent sets: $\varnothing$ has weight 0, $\{v\}$ has weight 1, and $\{w\}$ has weight 2. By difference of cubes,
$I_{(G,\omega)}(t)=\frac{1-t^{3}}{1-t}.$
Then by repeated difference of squares,
$\frac{1}{1-t}=(1+t)(1+t^{2})(1+t^{4})(1+t^{8})\dots=\prod_{j\geq 0}(1+t^{2^{j}}),$
and similarly,
$1-t^{3}=\frac{1}{1+t^{3}}\cdot\frac{1}{1+t^{6}}\cdot\frac{1}{1+t^{12}}\cdot\frac{1}{1+t^{24}}\dots=\prod_{j\geq 0}\frac{1}{1+t^{3\cdot 2^{j}}},$
which gives
$I_{(G,\omega)}(t)=\prod_{j\geq 0}(1+t^{2^{j}})\cdot\prod_{j\geq 0}\frac{1}{1+t^{3\cdot 2^{j}}}.$
The other subsets of $V(G)$ are $\{v\},$$\{w\},$ and $\varnothing$.∅ . In these cases, we get:
$I_{(G|_{\{v\}},\omega)}(t)$ $=1+t,$
$I_{(G|_{\{w\}},\omega)}(t)$ $=1+t^{2},$
$I_{(G|_{\varnothing},\omega)}(t)$ $=1.$
Putting this together, Theorem 1 gives
$\overline{X}_{(G,\omega)}(t)=\prod_{j\geq 0}(1+\overline{p}_{2^{j}})\cdot\prod_{j\geq 0}\frac{1}{1+\overline{p}_{3\cdot 2^{j}}}-(1+\overline{p}_{1})-(1+\overline{p}_{2})+1.$
Note that in this case, the superscript $2^{j}$ refers to the partition with one part whose size is $2^{j}$ rather than the partition with $j$ parts of size 2. It follows that all $\overline{p}$-coefficients are ±$\pm 1$ or 0 in this case. In particular,
$[\overline{p}_{\lambda}]\overline{X}_{(G,\omega)}=(-1)^{\#\textnormal{ of parts of }\lambda\textnormal{ that are multiples of 3}}$
provided that $\lambda$ satisfies all of the following properties:
  • $|\lambda|\geq 3.$
  • All parts are of the form $2^{j}$ or $3\cdot 2^{j}$ for some $j\geq 0.$
  • No parts of the form $2^{j}$ are repeated.
All other coefficients are 0. To illustrate this, we can expand out the first few terms to get
$\overline{X}_{(G,\omega)}(t)$ $=(1+\overline{p}_{1})(1+\overline{p}_{2})(1+\overline{p}_{4})\dots\frac{1}{1+\overline{p}_{3}}\cdot\frac{1}{1+\overline{p}_{6}}\cdot\frac{1}{1+\overline{p}_{12}}\cdot\dots-(1+\overline{p}_{1}+\overline{p}_{2})$
$=(1+\overline{p}_{1})(1+\overline{p}_{2})(1+\overline{p}_{4})\dots(1-\overline{p}_{3}+\overline{p}_{33}-\overline{p}_{333}+\dots)(1-\overline{p}_{6}+\overline{p}_{66}-+\dots)\dots-(1+\overline{p}_{1}+\overline{p}_{2})$
$=(-\overline{p}_{3}+\overline{p}_{21})+(\overline{p}_{4}-\overline{p}_{31})+(\overline{p}_{41}-\overline{p}_{32})+(-\overline{p}_{6}+\overline{p}_{42}+\overline{p}_{33}-\overline{p}_{321})$
$\ \ \ +(-\overline{p}_{61}+\overline{p}_{421}+\overline{p}_{331})+(\overline{p}_{8}-\overline{p}_{62}+\overline{p}_{332})+(\overline{p}_{81}+\overline{p}_{63}-\overline{p}_{621}-\overline{p}_{333}+\overline{p}_{3321})+\dots$
We can see that these terms match the ones from §1.
For the seemingly simpler case of a single unweighted edge [Uncaptioned image], there does not seem to be such a simple formula for the coefficients:
Example 9.
[…]
$I_{G}(t)=1+2t,$
since $\{v\}$ and $\{w\}$ are stable sets of size 1 and $\varnothing$ is a stable set of size 2. In this case the factorization begins
$1+2t=(1+t)^{2}\cdot\frac{1}{1+t^{2}}\cdot(1+t^{3})^{2}\cdot\frac{1}{(1+t^{4})^{4}}\cdot(1+t^{5})^{6}\dots.$
The other stable sets are $\{v\},$$\{w\}$, and $\varnothing$,∅ , which have independence polynomials
$I_{G|_{\{v\}}}(t)=I_{G|_{\{w\}}}(t)=1+t,\ \ \ \ I_{G|_{\varnothing}}(t)=1.$
Putting this together, Theorem 1 gives
$\overline{X}_{G}(t)$ $=(1+\overline{p}_{1})^{2}\cdot\frac{1}{1+\overline{p}_{2}}\cdot(1+\overline{p}_{3})^{2}\cdot\frac{1}{(1+\overline{p}_{4})^{4}}\cdot(1+\overline{p}_{5})^{6}\dots-(1+\overline{p}_{1})-(1+\overline{p}_{1})+1$
$=(1+2\overline{p}_{1}+\overline{p}_{11})(1-\overline{p}_{2}+\overline{p}_{22}-\dots)(1+2\overline{p}_{3}+\overline{p}_{33})(1-4\overline{p}_{4}+10\overline{p}_{44}-20\overline{p}_{444}+\dots)\dots-(1+2\overline{p}_{1})$
$=(-\overline{p}_{2}+\overline{p}_{11})+(2\overline{p}_{3}-2\overline{p}_{21})+(-4\overline{p}_{4}+4\overline{p}_{31}+\overline{p}_{22}-\overline{p}_{211})$
$\ \ \ +(6\overline{p}_{5}-8\overline{p}_{41}-3\overline{p}_{32}+2\overline{p}_{311}+2\overline{p}_{221})$
$\ \ \ +(-9\overline{p}_{6}+12\overline{p}_{51}+4\overline{p}_{42}-4\overline{p}_{411}+\overline{p}_{33}-4\overline{p}_{321}-\overline{p}_{222}+\overline{p}_{2211})\dots,$
again matching the expansion from §1. The sequence of exponents $2,-1,2,-4,6,\dots$ is sequence A038067 in the Online Encyclopedia of Integer Sequences (OEIS) [oeis].
4 Proofs
In this section we will prove the Theorem 1, Corollary 2, Proposition 3, and Theorem 4.
Proof of Theorem 1.
Our key lemma is an expansion formula for Stanley’s $Y_{G}$ function (introduced in [stanley1998graph]), where $Y_{(G,\omega)}$ is defined to be the same as $\overline{X}_{(G,\omega)}$ except that it includes monomials corresponding to colorings where some vertices are not assigned any colors:
$Y_{(G,\omega)}:=\sum_{\kappa}\prod_{v\in V(G)}\left(\prod_{i\in\kappa(v)}x_{i}\right)^{\omega(v)},$
where the sum ranges over all functions $\kappa:V(G)\to 2^{n}$ such that $\kappa(v)\cap\kappa(w)=\varnothing$ whenever $vw\in E(G).$
Lemma 10.
$Y_{(G,\omega)}$ factors as
$Y_{(G,\omega)}=\prod_{k\geq 1}(1+\overline{p}_{k})^{a_{V(G)}(k)}.$
Proof.
Since $Y_{(G,\omega)}$ allows vertices to not receive any colors, we can think of assigning each color to a subset of the vertices independently of all other colors. Specifically, for each color, we can independently choose a subset of $V(G)$ which will receive that color, and the only restriction is that the subset be an independent set. Thus, the generating series for the ways to choose which vertices receive color $i$ is given by the independence polynomial $I_{(G,\omega)}(x_{i})$. Since different colors can be assigned independently, we can compute $Y_{(G,\omega)}$ by simply multiplying over all colors, so
$Y_{(G,\omega)}=\prod_{i\geq 1}I_{(G,\omega)}(x_{i}).$
Then by definition of $a_{V(G)}(k)$, we have
$I_{(G,\omega)}(x_{i})=\prod_{k\geq 1}(1+x_{i}^{k})^{a_{V(G)}(k)}.$
Now note that $\overline{p}_{k}$ represents the sum of all monomials corresponding to ways to assign a nonempty set of colors to a single weight $k$ vertex, so $1+\overline{p}_{k}$ represents the sum of monomials corresponding to all ways to assign a set of colors to the weight $k$ vertex, including the empty set. To choose such a coloring, we can independently choose whether or not to use each color. If color $i$ is used, then we get an $x_{i}^{k}$ in the monomial, and otherwise the power of $x_{i}$ is 0. Since we can independently choose an exponent of 0 or $k$ for each variable,
$1+\overline{p}_{k}=\prod_{i\geq 1}(1+x_{i}^{k}).$
It follows that
$Y_{(G,\omega)}=\prod_{i\geq 1}I_{(G,\omega)}(x_{i})=\prod_{i,k\geq 1}(1+x_{i}^{k})^{a_{V(G)}(k)}=\prod_{k\geq 1}(1+\overline{p}_{k})^{a_{V(G)}(k)},$
as claimed. ∎
Theorem 1 then follows from the principle of inclusion-exclusion. The difference between $\overline{X}_{(G,\omega)}$ and $Y_{(G,\omega)}$ is that $\overline{X}_{(G,\omega)}$ requires all vertices to receive a nonempty set of colors. Thus, by inclusion-exclusion, we can compute $\overline{X}_{(G,\omega)}$ by adding the monomials all possible colorings without that restriction, then subtracting the monomials where a particular vertex receives no colors, then adding back the ones where a particular pair of vertices receive no colors, and so on. It follows that
$\overline{X}_{(G,\omega)}=\sum_{W\subseteq V(G)}Y_{(G,\omega)}.$
Plugging in the formula from Lemma 10 proves Theorem 1. ∎
The formula in Corollary 2 is now fairly immediate:
Proof of Corollary 2.
Let $\lambda=\ell^{i_{\ell}}\dots 2^{i_{2}}1^{i_{1}}.$ First, we note that
$\overline{p}_{\lambda}=\overline{p}_{\lambda_{1}}\dots\overline{p}_{\lambda_{\ell(\lambda)}}=\overline{p}_{\ell}^{i_{\ell}}\dots\overline{p}_{2}^{i_{2}}\overline{p}_{1}^{i_{1}}.$
To see this, recall that $\overline{p}_{\lambda}$ is the sum of all monomials corresponding to colorings of the vertices of the edgeless graph $\overline{K_{\lambda}}$ with vertex weights $\lambda_{1},\lambda_{2},\dots,\lambda_{\ell(\lambda)}$, or equivalently, with $i_{k}$ vertices of weight $k$ for each $k=1,2,\dots,\ell.$ Since the graph has no edges, the set of colors assigned to a vertex has no bearing on how we can color the other vertices, so we can choose the coloring of each vertex independently. Thus, $\overline{p}_{\lambda}$ is the product over all vertices of the sum of all possible monomials for ways to color that vertex, which is just $\overline{p}_{k}$ for a vertex of weight $k.$
Thus, get a $\overline{p}_{\lambda}$ term from $\prod_{k\geq 1}(1+\overline{p}_{k})^{a_{W}(k)},$ we need to choose a $\overline{p}_{k}^{i_{k}}$ term from $(1+\overline{p}_{k})^{a_{W}(k)}$ for each $k=1,2,\dots,\ell.$ By the binomial theorem,
$(1+\overline{p}_{k})^{a_{W}(k)}=\sum_{i\geq 0}\binom{a_{W}(k)}{i}\overline{p}_{k}^{i},$
so the coefficient of $\overline{p}_{\lambda}$ in the product $\prod_{k\geq 1}(1+\overline{p}_{k})^{a_{W}(k)}$ is just the product of these coefficients:
$[\overline{p}_{\lambda}]\prod_{k\geq 1}(1+\overline{p}_{k})^{a_{W}(k)}=\prod_{k=1}^{\ell}\binom{a_{W}(k)}{i_{k}}\overline{p}_{k}^{i}.$
Summing over all $W\subseteq V(G)$ completes the proof. ∎
Next, we prove the recursive formula in Proposition 3:
Proof of Proposition 3.
Taking logarithms of both sides of the equation
$(1+t)^{a_{W}(1)}(1+t^{2})^{a_{W}(2)}(1+t^{3})^{a_{W}(3)}\dots=I_{(G|_{W},\omega)}(t)$
gives
$\sum_{j\geq 1}a_{W}(j)\log(1+t^{j})=\log(I_{G|_{W},\omega}(t)).$
Expanding the left side using the power series $\log(1+x)=\sum_{i\geq 1}\frac{(-1)^{i+1}x^{i}}{i},$ we get
$\sum_{j\geq 1}a_{W}(j)\sum_{i\geq 1}\frac{(-1)^{i+1}t^{ij}}{i}=\log(I_{G|_{W},\omega}(t)).$
Since $j=k$ is the largest value of $j$ on the left side that contributes a $t^{k}$ term, we can compute $a_{W}(k)$ recursively from the values $a_{W}(j)$ for $j<k$ by making the coefficient of $t^{k}$ on the left side match the coefficient $[t^{k}]\log(I_{G|_{W},\omega}(t))$ on the right side. The values $j<k$ that contribute a $t^{k}$ term on the left are precisely the values where $j$ is some divisor $d$ of $k$, in which case $i=k/d.$ The contribution from the $j=d$ term is then
$a_{W}(d)\frac{(-1)^{k/d+1}}{k/d}=\frac{(-1)^{k/d}d\cdot a_{W}(d)}{k}.$
Summing over all values $d\mid k$ and then subtracting the resulting sum from the right side coefficient $[t^{k}]\log(I_{G|_{W},\omega}(t))$ gives the formula in Proposition 3. ∎
Finally, we prove that these coefficients are all integers:
Proof of Theorem 4.
Consider the $\overline{\widetilde{m}}$-expansion of $\overline{p}_{\lambda}=\overline{X}_{\overline{K_{\lambda}}}.$ Since $\overline{K_{\lambda}}$ has no edges, all subsets of its vertices are independent sets. One possible stable set cover is the one where each vertex is in a stable set by itself, in which case $\lambda(C)=\lambda,$ so the expansion of $\overline{p}_{\lambda}$ has an $\overline{\widetilde{m}}_{\lambda}$ term. For every other term $\overline{\widetilde{m}}_{\mu}$, we must have either $|\mu|>|\lambda|$ (if some vertices are covered by multiple stable sets in $C$) or else $\mu$ is formed by taking $\lambda$ and combining some of its parts (if each vertex is only in one stable set but some stable sets contain multiple vertices), which is equivalent to saying that $\lambda$ is a refinement of $\mu$. Thus, we have $[\overline{\widetilde{m}}_{\lambda}]\overline{p}_{\lambda}=1$ and $[\overline{\widetilde{m}}_{\mu}]\overline{p}_{\mu}=0$ whenever $|\mu|<|\lambda|$ or $|\mu|=|\lambda|$ but $\lambda$ is not a refinement of $\mu.$ Now, order the set of all partitions in such a way that partitions of smaller numbers come first and $\lambda$ comes before $\mu$ whenever $\lambda$ is a refinement of $\mu$, which can be done by ordering the partitions of each size lexicographically. Then the transition matrix from the $\overline{p}$-basis to the $\overline{\widetilde{m}}$-basis is upper triangular with 1’s along the diagonal and all integer entries. It follows that the inverse matrix which transitions from the $\overline{\widetilde{m}}$-basis to the $\overline{p}$-basis is also upper triangular with 1’s on the diagonal and all integer entries. Thus, each $\overline{\widetilde{m}}_{\mu}$ is a linear combination of $\overline{p}_{\lambda}$ terms with integer coefficients. Since $\overline{X}_{(G,\omega)}$ is an integral linear combination of $\overline{\widetilde{m}}_{\mu}$ terms for every weighted graph $(G,\omega)$, it follows that $\overline{X}_{(G,\omega)}$ is an integral linear combination of $\overline{p}_{\lambda}$ terms, as claimed. ∎
Acknowledgements
The author thanks Oliver Pechenik for suggesting the problem, and she thanks Oliver Pechenik, Karen Yeats, and Sophie Spirkl for providing helpful comments. She was partially supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) grant RGPIN-2022-03093.
[…]